OpenStudy (anonymous):

To the nearest tenth, what is the area of the shaded segment when JA = 8 ft?

jimthompson5910 (jim_thompson5910):

what is the area of the full circle?

OpenStudy (anonymous):

how do you find that out

jimthompson5910 (jim_thompson5910):

using the formula A = pi*r^2

jimthompson5910 (jim_thompson5910):

r = radius

OpenStudy (anonymous):

i got 188.4 idk if its right

jimthompson5910 (jim_thompson5910):

what is the radius

OpenStudy (anonymous):

i honestly dont know how to do this problem at all

jimthompson5910 (jim_thompson5910):

that's ok, can you tell me what the radius is?

OpenStudy (anonymous):

8 i think

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = pi*8^2 A = ???

OpenStudy (anonymous):

25.12

jimthompson5910 (jim_thompson5910):

8^2 means 8 squared so 8^2 = 8*8 = ??

OpenStudy (anonymous):

64

jimthompson5910 (jim_thompson5910):

multiply that with pi or 3.14 roughly

OpenStudy (anonymous):

200.96

jimthompson5910 (jim_thompson5910):

that's the area of the circle (roughly)

jimthompson5910 (jim_thompson5910):

now what is the area of the pie shaped wedge? ie this shaded region